# Higher-order homogenised riblet boundary conditions

### v1

@@ 2.2.1 Order zero
At outer order 0 we need to solve the Laplace equation for $\phi_{0}(y,z)$ with the boundary condition (2.6) at the outer wall and $\phi_{0}(y,0)=0$ at the flat inner wall $z=0$, to which the periodic wall reduces for $\varepsilon=0$. From the result we can extract $\phi_{0,z}(y,0)$ (which also includes information on all of its $y$-derivatives) for later use.
@@ 2.2.2 First order
where $y_{0}$ is to be treated as a parameter as far as the Laplace equation for $\varPhi_{1}$ is concerned, and symbol $\sim$ is here precisely defined to mean that the difference between its l.h.s. and r.h.s. must tend to zero when $Z\rightarrow\infty$. This definition may imply more than one actual equation, since the l.h.s. and r.h.s. are asymptotically polynomials in $Z$, and each coefficient of these polynomials must individually match for the difference to tend to zero. Notice also that the second term of (2.7), $Y\phi_{0,y}(y_{0},0)$, has been omitted from (2.8) because $\phi_{0}(y,0)=0$ for all $y$, and its derivatives are thus also zero. Asymptotic condition (2.8) can be split into the two equations
@@
and, within a similar error O$(\varepsilon^{2})$, replace $\varepsilon\phi_{0,z}(y,0)$ by $\varepsilon\phi_{z}(y,0)$ (similarly, in a sense, to what in Aerodynamics is called “interactive boundary layer theory”, e.g. Cebeci (1999)). There results a first-order equivalent boundary condition for the total $\phi$ without any more reference to its series expansion:
@@ 2.2.3 Second order
where $\phi_{0,yy}$ has been omitted because $\phi_{0}(y,0)$ is identically zero, and $\phi_{0,zz}$ has also been omitted for the same reason, because from the Laplace equation it equals $-\phi_{0,yy}$. We notice also that, from the $y$-derivative of (2.11), $\phi_{1,y}(y_{0},0)$ can be replaced by $h_{1}\phi_{0,yz}(y_{0},0)$.
@@
The new dimensionless coefficient $h_{2}$ will be called “second-order protrusion height” for naming consistency, but it should be evident that $s^{*2}h_{2}$ is an area and not a height. Also, it can be noted that if $\overline{\varPhi_{21}}$ happens to be an odd function of $Y$, a limit such as (2.23), which must then equal its opposite under the transformation $Y\leftarrow-Y$, is necessarily zero. More about this symmetry at the end of §2.2.5.
@@ 2.2.5 Compound equivalent boundary condition
Just as (2.14) was at first order, this single implicit $N^{th}$-order boundary condition is free of any reference to the outer series expansion, and may be preferred to (2.32) for computational purposes. One can also observe that if the wall shape $Z=F(Y)$ happens to be left-right symmetric, i.e. invariant under the transformation $Y\leftarrow-Y$, only even $y$-derivatives have the same symmetry. Odd $y$-derivatives are forbidden to appear in (2.32)–(2.35), and the corresponding higher-order protrusion heights must automatically come out as zero. Notice that this does not imply that the corresponding inner solutions are zero, only that they shall have the appropriate symmetry. For instance, when riblets are symmetric $\overline{\varPhi_{21}}$ of (2.22) is an odd function of $Y$, because so is $\overline{\varPhi_{11}}_{,Y}$, and its mean value, which when $Z\rightarrow\infty$ provides $h_{2}$, must be zero.
@@ 3 Longitudinal Stokes flow past riblets
where the, typically negative, pressure gradient $p^{*}_{x^{*}}$ is of an order of magnitude dictated by the outer scale, and for this reason was neglected by Bechert & Bartenwerfer (1989) and by Luchini et al. (1991). When the pressure gradient is neglected, the longitudinal velocity $u^{*}$ behaves exactly like the quantity $\phi^{*}$ of §2.
At higher than first order in $\varepsilon$, the pressure gradient can no longer be neglected and, as will be seen, it provides a correction even for constant outer conditions. For the purpose of finding this correction we shall assume the pressure gradient $p^{*}_{x^{*}}/\mu^{*}$ to be proportional to $u^{*}/L^{*2}$, so that (3.2) is balanced when both sides are expressed in outer units. In other words, the outer dimensionless pressure gradient $p_{x}=L^{*2}p^{*}_{x^{*}}/\mu^{*}u_{R}^{*}$ (where $u_{R}^{*}$ is a reference velocity that will be specified later but for the time being can be arbitrary) is of order unity. Then the dimensionless form of (3.2) on the scale of the period $s^{*}$ (i.e. in inner units) is
@@ 3.1 Wavenumber domain (constant outer conditions)
[…]
@@
which includes the second-order effect of the pressure gradient. Again no higher powers of $\varepsilon$ are involved, and (3.6) is exponentially accurate, when outer conditions are constant and the solution is perfectly periodic.
@@ 3.2 Physical domain (slowly varying outer conditions)
To remind what a Cauchy series is, it is in fact a Taylor series in a neighbourhood of a generic wall point $(y_{0},0)$, where $z$-derivatives higher that the first have been replaced using the differential equation differentiated zero or more times in $z$:
@@
This device was originally invented by Cauchy in order to express the general solution of an initial-value problem in time, but here it will be used to reformulate our boundary-value problem as an initial-value problem in $z$. Whereas for an elliptic equation the initial-value problem is notoriously ill-posed, the trick here will be to formally apply the Cauchy series to boundary values obtained from the two-sided boundary-value problem itself. This will allow us to reformulate the inner problem as an equivalent boundary condition, without (or before) actually solving the outer problem.
@@ 4 Transverse 2D Stokes problem
This is a $4^{th}$-order elliptic equation requiring 4 boundary conditions, 2 at each side of a strip-like domain in $z^{*}$. At the solid riblet wall we have $\partial\psi^{*}/\partial z^{*}=\partial\psi^{*}/\partial y^{*}=0$, which is the same as $\partial\psi^{*}/\partial n^{*}=0$ (normal to the riblet wall, i.e. no slip) and $\partial\psi^{*}/\partial s^{*}=0$ (parallel to the riblet wall, i.e. impermeable). The latter can be integrated to $\psi^{*}=\psi^{*}_{\text{w}}$, an arbitrary constant which on one of the two walls can without loss of generality be taken as $\psi^{*}_{\text{w}}=0$.
@@
$\left[\begin{matrix}s^{*}\psi^{*}_{z^{*}}(y,0)\\ \psi^{*}(y^{*},0)\end{matrix}\right]=\underbrace{\left[\begin{matrix}a_{1}&b_{2}\\ c_{2}&d_{3}\\ \end{matrix}\right]}_{\text{dimensionless}}\left[\begin{matrix}s^{*2}\psi^{*}_{z^{*}z^{*}}(y^{*},0)\\ s^{*3}\psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},0)\end{matrix}\right],$
@@ 4.2.1 Order zero
With $\psi(y,0)$, $\psi_{z}(y,0)$, $\psi_{zz}(y,0)$, and $\psi_{zzz}(y,0)$ known, we can formally reconstruct the outer solution by Cauchy series (actually by Taylor series until, at fourth order, we begin to see the effect of the biharmonic differential equation in eliminating fourth and higher $z$-derivatives):
$\begin{split}&\psi(y,z)=\psi(y_{0},0)+(y-y_{0})\psi_{y}(y_{0},0)+z\psi_{z}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{2}}{2}\psi_{yy}(y_{0},0)+(y-y_{0})z\psi_{yz}(y_{0},0)+\frac{z^{2}}{2}\psi_{zz}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{3}}{6}\psi_{yyy}(y_{0},0)+\frac{(y-y_{0})^{2}z}{2}\psi_{yyz}(y_{0},0)+\frac{(y-y_{0})z^{2}}{2}\psi_{yzz}(y_{0},0)+\frac{z^{3}}{6}\psi_{zzz}(y_{0},0)+{}\\ &\frac{z^{4}}{24}\left[-\psi_{yyyy}(y_{0},0)-2\psi_{yyzz}(y_{0},0)\right]+\cdots\end{split}$ (4.10)
@@
[…]
The cancellations are due to $\psi_{0}$ and $\psi_{0,z}$ being zero, together with their $y$-derivatives, on the flat wall that corresponds to $\varepsilon=0$. This is the sequence that must be matched, term by term according to powers of $\varepsilon$, to the inner expansion (4.8).
@@ 4.2.3 Second order
where $\overline{\varPsi_{21}}(Y,Z)$ was defined in (4.14), and $\overline{\varPsi_{31}}(Y,Z)$ is a periodic function of $Y$ obeying the equation
@@
The second term of (4.20) proportional to $Z^{3}/6$ can be matched by a normalized contribution $\overline{\varPsi_{32}}$ which obeys the biharmonic equation with boundary conditions
@@ 4.2.4 Third order
$\begin{split}\psi_{3}(y_{0},0)=&\lim_{Z\rightarrow\infty}\left[\varPsi_{3}-\frac{YZ^{2}}{2}\psi_{0,yzz}-\frac{Z^{3}}{6}\psi_{0,zzz}-\frac{Z^{2}}{2}\psi_{1,zz}-Y\psi_{2,y}-Z\psi_{2,z}\right]=\\ &c_{3}\psi_{0,yzz}+c_{2}\psi_{1,yzz}+d_{3}\psi_{0,zzz},\end{split}$ (4.26)
@@
Constant $d_{3}$, just as well as its previous sister constants, equals $d_{3}$ of §4.1.
@@
where $\overline{\varPsi_{21}}$ and $\overline{\varPsi_{31}}$ were defined earlier, and $\overline{\varPsi_{41}}(Y,Z)$ is a periodic function of $Y$ obeying the equation
@@
the latter two coming from the $-Z^{4}/12$ behaviour at $Z\rightarrow\infty$ expected of $\overline{\varPsi_{41}}$ in (4.29).
@@
where $\overline{\varPsi_{32}}$ was defined earlier, and $\overline{\varPsi_{42}}(Y,Z)$ is a periodic function of $Y$ obeying the equation
@@
$\begin{split}\varPsi_{4}=&\left[\overline{\varPsi_{41}}(Y,Z)+Y\overline{\varPsi_{31}}(Y,Z)+\frac{Y^{2}}{2}\overline{\varPsi_{21}}(Y,Z)\right]\psi_{0,yyzz}+\left[\overline{\varPsi_{42}}(Y,Z)+Y\overline{\varPsi_{32}}(Y,Z)\right]\psi_{0,yzzz}+{}\\ &\left[\overline{\varPsi_{31}}(Y,Z)+Y\overline{\varPsi_{21}}(Y,Z)\right]\psi_{1,yzz}+\overline{\varPsi_{32}}(Y,Z)\psi_{1,zzz}+\overline{\varPsi_{21}}(Y,Z)\psi_{2,zz}\end{split}$ (4.30)
@@
Equation (4.31), together with (4.26), closes the problem for the third-order outer solution $\psi_{3}$. Owing to special symmetries of the biharmonic equation, it also turns out that $b_{3}=-c_{3}$.
@@ 4.2.5 Higher orders and equivalent boundary condition
Each $n^{th}$ matching condition will contain two new contributions, one proportional to $\partial^{n}\psi_{0}/\partial y^{n-2}\partial z^{2}$ and one proportional to $\partial^{n}\psi_{0}/\partial y^{n-3}\partial z^{3}$, in addition to a sequence of other terms which mimic the $(n-1)^{th}$ matching condition with $\psi_{i}$ replaced by $\psi_{i+1}$. The two new contributions will produce two independent normalized inner solutions $\overline{\varPsi_{n1}}(Y,Z)$ and $\overline{\varPsi_{n2}}(Y,Z)$, and eventually produce constant numerical coefficients $a_{n-1}$, $b_{n-1}$, $c_{n}$, and $d_{n}$ for the $n^{th}$ outer boundary condition
@@
Also, as was true in §2.2.5, when the geometry is left-right symmetric (4.33)–(4.35) must enforce the same symmetry, which means that no odd $y$-derivatives of $\psi$ can appear. The $a_{i},b_{i},c_{i},d_{i}$ coefficients corresponding to odd derivatives will automatically turn out to be zero in this case.

### v2

@@ 2.2.1 Order zero
At outer order 0 we need to solve the Laplace equation for $\phi_{0}(y,z)$ with the boundary condition (2.6) at the outer wall and $\phi_{0}(y,0)=0$ at the flat inner wall $z=0$, to which the periodic wall reduces for $\varepsilon=0$. From the result we can extract $\phi_{0,z}(y,0)$ (which also includes information about all of its $y$-derivatives) for later use.
@@ 2.2.2 First order
where $y_{0}$ is to be treated as a parameter as far as the Laplace equation for $\varPhi_{1}$ is concerned, and symbol $\sim{}$ (“behaves as”) is here precisely defined to mean that the difference between its l.h.s. and r.h.s. must tend to zero when $Z\rightarrow\infty$. This definition may imply more than one actual equation, since the l.h.s. and r.h.s. are asymptotically polynomials in $Z$, and each coefficient of these polynomials must individually match for the difference to tend to zero. Notice also that the second term of (2.7), $Y\phi_{0,y}(y_{0},0)$, has been omitted from (2.8) because $\phi_{0}(y,0)=0$ for all $y$, and its derivatives are thus also zero. With $\sim{}$ given the above meaning, asymptotic condition (2.8) is equivalent to the two equations
@@
[…]
@@ 2.2.3 Second order
where $\phi_{0,yy}$ has been omitted because $\phi_{0}(y,0)$ is identically zero, and $\phi_{0,zz}$ has also been omitted for the same reason, because from the Laplace equation it equals $\phi_{0,yy}$. We notice also that, from the $y$-derivative of (2.11), $\phi_{1,y}(y_{0},0)$ can be replaced by $h_{1}\phi_{0,yz}(y_{0},0)$.
@@
The new dimensionless coefficient $h_{2}$ can be called “second-order protrusion coefficient” (or “second-order protrusion height”, but we refrain from this denomination because $s^{*2}h_{2}$ has physical dimensions of the square of a height, and subsequent $n^{\text{th}}$-order coefficients will be the $n^{\text{th}}$ power of a height). Also, it can be noted that if $\overline{\varPhi_{21}}$ happens to be an odd function of $Y$, a limit such as (2.23), which must then equal its opposite under the transformation $Y\leftarrow-Y$, is necessarily zero. More about this symmetry at the end of §2.2.5.
@@ 2.2.5 Compound equivalent boundary condition
Just as (2.14) was at first order, this single implicit $N^{\text{th}}$-order boundary condition is free of any reference to the outer series expansion, and may be preferred to (2.32) for computational purposes. One can also observe that if the wall shape $Z=F(Y)$ happens to be left-right symmetric, i.e. invariant under the mirror transformation $Y\leftarrow-Y$, only even $y$-derivatives have the same symmetry and are allowed in (2.32)–(2.35). Odd $y$-derivatives are forbidden by symmetry, and the corresponding higher-order protrusion coefficients must automatically come out as zero. Notice that this does not imply that the corresponding inner solutions are zero, only that they shall have the appropriate symmetry for these coefficients to vanish. For instance, when riblets are mirror symmetric, $\overline{\varPhi_{21}}$ of (2.22) is an odd function of $Y$, because so is $\overline{\varPhi_{11}}_{,Y}$, and thus its mean value, which when $Z\rightarrow\infty$ provides $h_{2}$, is zero.
@@ 3 Longitudinal Stokes flow past riblets
[…]
At higher than first order in $\varepsilon$, the pressure gradient can no longer be neglected and, as will be seen, it provides a correction even for constant outer conditions. Exactly where in the expansion this correction appears will depend on the particular distinguished limit adopted. For the purpose of a flow where the pressure gradient retains the same order of magnitude it had on a flat wall (i.e. where the ratio of the pressure gradient to outer scales is kept constant while $\varepsilon\rightarrow 0$), we shall assume the pressure gradient $p^{*}_{x^{*}}/\mu^{*}$ to be proportional to $u^{*}/L^{*2}$, so that (3.2) is balanced when both sides are expressed in outer units. In other words, the outer dimensionless pressure gradient $p_{x}=L^{*2}p^{*}_{x^{*}}/\mu^{*}u_{R}^{*}$ (where $u_{R}^{*}$ is a reference velocity that will be specified later but for the time being can be arbitrary) is kept constant and of order unity. Then the dimensionless form of (3.2) on the scale of the period $s^{*}$ (i.e. in inner units) is
@@ 3.1 Wavenumber domain (constant outer conditions)
by linearity, the two constants $h_{1}$ and $h^{(p_{x})}_{2}$ can be determined through once applying a suitable non-zero boundary condition on $\hat{U}_{0,Z}$ at $Z=L$ with $p_{x}=0$ (and this shows that $h_{1}$ is precisely the same protrusion height as defined in §2), and then the boundary condition $\hat{U}_{0,Z}=-\varepsilon^{2}p_{x}L$ at $Z=L$, to make $b_{0}=0$ with $p_{x}\neq 0$. Since $b_{0}=0,$ and $\hat{U}_{0}(L)=\cancel{b_{0}(h_{1}+L)}+\varepsilon^{2}p_{x}(h^{(p_{x})}_{2}+L^{2}/2),$ one can obtain
@@
which includes the second-order effect of the pressure gradient. Again no higher powers of $\varepsilon$ are involved, and (3.6) is exponentially accurate, when outer conditions are constant and the inner solution is perfectly periodic.
@@ 3.2 Physical domain (slowly varying outer conditions)
To remind what a Cauchy series is, it is in fact a Taylor series in a neighbourhood of a generic wall point $(y_{0},0)$, where $z$-derivatives higher that the first have been replaced using the differential equation, differentiated zero or more times in $z$:
@@
This device was originally invented by Cauchy in order to express the general solution of an initial-value problem in time, but here it will be used to reformulate our boundary-value problem as an initial-value problem in $z$. Whereas for an elliptic equation the initial-value problem is notoriously ill-posed, the trick here will be to formally apply the Cauchy series to initial values obtained from the two-sided boundary-value problem itself. This will allow us to reformulate the inner problem as an equivalent boundary condition, without (or before) actually solving the outer problem.
@@ 4 Transverse 2D Stokes problem
This is a $4^{\text{th}}$-order elliptic equation requiring 4 boundary conditions, 2 at each side of a strip-like domain in $z^{*}$. At the solid riblet wall we have $\partial\psi^{*}/\partial z^{*}=\partial\psi^{*}/\partial y^{*}=0$, which is the same as $\partial\psi^{*}/\partial n^{*}=0$ (derivative normal to the riblet wall provides the tangential velocity, i.e. no slip) and $\partial\psi^{*}/\partial s^{*}=0$ (derivative parallel to the riblet wall provides the normal velocity, i.e. impermeability). $\partial\psi^{*}/\partial s^{*}=0$ can be integrated to $\psi^{*}=\psi^{*}_{\text{w}}$, an arbitrary constant which on one of the two walls can without loss of generality be taken as $\psi^{*}_{\text{w}}=0$.
@@
$\left[\begin{matrix}s^{*}\psi^{*}_{z^{*}}(y,0)\\ \psi^{*}(y^{*},0)\end{matrix}\right]=\underbrace{\left[\begin{matrix}a_{1}&b_{2}\\ c_{2}&d_{3}\\ \end{matrix}\right]}_{\text{dimensionless}}\left[\begin{matrix}s^{*2}\psi^{*}_{z^{*}z^{*}}(y^{*},0)\\ s^{*3}\psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},0)\end{matrix}\right],$ (4.4)
@@ 4.2.1 Order zero
With $\psi(y,0)$, $\psi_{z}(y,0)$, $\psi_{zz}(y,0)$, and $\psi_{zzz}(y,0)$ all known, we can formally reconstruct the outer solution by Cauchy series (actually by Taylor series until, at fourth order, we begin to see the effect of the biharmonic differential equation in eliminating fourth and higher $z$-derivatives):
$\begin{split}&\psi(y,z)=\psi(y_{0},0)+(y-y_{0})\psi_{y}(y_{0},0)+z\psi_{z}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{2}}{2}\psi_{yy}(y_{0},0)+(y-y_{0})z\psi_{yz}(y_{0},0)+\frac{z^{2}}{2}\psi_{zz}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{3}}{6}\psi_{yyy}(y_{0},0)+\frac{(y-y_{0})^{2}z}{2}\psi_{yyz}(y_{0},0)+\frac{(y-y_{0})z^{2}}{2}\psi_{yzz}(y_{0},0)+{}\\ &\frac{z^{3}}{6}\psi_{zzz}(y_{0},0)+\frac{z^{4}}{24}\left[-\psi_{yyyy}(y_{0},0)-2\psi_{yyzz}(y_{0},0)\right]+\cdots\end{split}$ (4.11)
@@
$\begin{split}&\psi(y,z)=\cancel{\psi_{0}(y_{0},0)}+\varepsilon\left(Y\cancel{\psi_{0,y}}+Z\cancel{\psi_{0,z}}+\psi_{1}(y_{0},0)\right)+{}\\ &\varepsilon^{2}\left(\frac{Y^{2}}{2}\cancel{\psi_{0,yy}}+YZ\cancel{\psi_{0,yz}}+\frac{Z^{2}}{2}\psi_{0,zz}+Y\psi_{1,y}+Z\psi_{1,z}+\psi_{2}(y_{0},0)\right)+{}\\ &\varepsilon^{3}\left(\frac{Y^{3}}{6}\cancel{\psi_{0,yyy}}+\frac{Y^{2}Z}{2}\cancel{\psi_{0,yyz}}+\frac{YZ^{2}}{2}\psi_{0,yzz}+\frac{Z^{3}}{6}\psi_{0,zzz}+{}\right.\\ &\left.\frac{Y^{2}}{2}\psi_{1,yy}+YZ\psi_{1,yz}+\frac{Z^{2}}{2}\psi_{1,zz}+Y\psi_{2,y}+Z\psi_{2,z}+\psi_{3}(y_{0},0)\right)+\cdots\end{split}$ (4.12)
The cancellations are due to $\psi_{0}$ and $\psi_{0,z}$ being zero, together with their $y$-derivatives, on the flat wall that corresponds to $\varepsilon=0$. This is the sequence that must be matched, term by term according to powers of $\varepsilon$, by the inner expansion (4.9).
@@ 4.2.3 Second order
where $\overline{\varPsi_{21}}(Y,Z)$ is the same as was defined in (4.15), and $\overline{\varPsi_{31}}(Y,Z)$ is a new periodic function of $Y$ obeying the equation
@@
The second term of (4.21) proportional to $Z^{3}/6$ can be matched by a normalized contribution [^1] Here and in what follows we shall adopt a convention where contributions are numbered by a first digit denoting order and a second sequential digit. $\overline{\varPsi_{32}}$ which obeys the biharmonic equation with boundary conditions
@@ 4.2.4 Third order
$\begin{split}\psi_{3}(y_{0},0)=&\lim_{Z\rightarrow\infty}\left[\varPsi_{3}-\frac{YZ^{2}}{2}\psi_{0,yzz}-\frac{Z^{3}}{6}\psi_{0,zzz}-\frac{Z^{2}}{2}\psi_{1,zz}-Y\psi_{2,y}-Z\psi_{2,z}\right]=\\ &c_{3}\psi_{0,yzz}+c_{2}\psi_{1,zz}+d_{3}\psi_{0,zzz},\end{split}$ (4.27)
@@
Constant $d_{3}$, just as well as its previous sister constants, coincides with $d_{3}$ of §4.1.
@@
where $\overline{\varPsi_{21}}$ and $\overline{\varPsi_{31}}$ were defined earlier, and $\overline{\varPsi_{41}}(Y,Z)$ is a new periodic function of $Y$ obeying the equation
@@
the latter two coming from the $Z^{4}/12$ behaviour at $Z\rightarrow\infty$ expected of $\overline{\varPsi_{41}}$ in (4.30).
@@
where $\overline{\varPsi_{32}}$ was defined earlier, and $\overline{\varPsi_{42}}(Y,Z)$ is yet another periodic function of $Y$ obeying the equation
@@
$\begin{split}\varPsi_{4}=&\left[\overline{\varPsi_{41}}(Y,Z)+Y\overline{\varPsi_{31}}(Y,Z)+\frac{Y^{2}}{2}\overline{\varPsi_{21}}(Y,Z)\right]\psi_{0,yyzz}+{}\\ &\left[\overline{\varPsi_{42}}(Y,Z)+Y\overline{\varPsi_{32}}(Y,Z)\right]\psi_{0,yzzz}+{}\\ &\left[\overline{\varPsi_{31}}(Y,Z)+Y\overline{\varPsi_{21}}(Y,Z)\right]\psi_{1,yzz}+\overline{\varPsi_{32}}(Y,Z)\psi_{1,zzz}+\overline{\varPsi_{21}}(Y,Z)\psi_{2,zz}\end{split}$ (4.31)
@@
Equation (4.32), together with (4.27), closes the problem for the third-order outer solution $\psi_{3}$. Owing to special symmetries of the biharmonic equation, it also turns out that $b_{3}=c_{3}$.
@@ 4.2.5 Higher orders and equivalent boundary condition
Each $n^{\text{th}}$ matching condition will contain two new contributions, one proportional to $\partial^{n}\psi_{0}/\partial y^{n-2}\partial z^{2}$ and one proportional to $\partial^{n}\psi_{0}/\partial y^{n-3}\partial z^{3}$, in addition to a sequence of other terms which mimic the $(n-1)^{\text{th}}$ matching condition with $\psi_{i}$ replaced by $\psi_{i+1}$. The two new contributions will produce two independent normalized inner solutions $\overline{\varPsi_{n1}}(Y,Z)$ and $\overline{\varPsi_{n2}}(Y,Z)$, and eventually produce constant numerical coefficients $a_{n-1}$, $b_{n-1}$, $c_{n}$, and $d_{n}$ for the $n^{\text{th}}$ outer boundary condition
@@
Also, as was true in §2.2.5, when the geometry is left-right symmetric (4.34)–(4.36) must enforce the same symmetry, which means that only even $y$-derivatives of $\psi$ can appear. The $a_{i},b_{i},c_{i},d_{i}$ coefficients corresponding to odd derivatives will automatically turn out to be zero in this case.